Macomb Corporation is a U.S. firm that invoices some of its exports in Japanese yen. If...

  1. Macomb Corporation is a U.S. firm that invoices some of its exports in Japanese yen. If it expects
    the yen to weaken, it could ____ to hedge the exchange rate risk on those exports.
    a. sell yen put options
    b. buy yen put options
    c. buy yen call options
    d. sell yen call options

with explanation please
Answer #1

Correct answer is option B - Buy Yen Put options

Reason:

Macomb Corp. is a seller which is expected to receive payments in Japanese.Yen. This international investment is exposed to currency volatility and exchange risk.

Example: Macomb sold/exported products worth 10000 Japanese Yen. This is expected to be received after 1 month.

Suppose, currently

$1 = 100 Japanese Yen.

So, Macomb is expected to get: 10000 / 100 Japanese Yen = $ 100

Now after 1 month exchange rate changes and Yen weaken

$1 = 110 Japanese Yen.

So, Macomb is expected to get: 10000 / 110 Japanese Yen = $ 90.909

Hence, there is a possibility of loss.

So to hedge buy Yen Put options. Put options give the choice to Macomb to Sell Yen at a pre defined price.

Hence, On one hand Macomb will receive Japanese Yen from sale and on other hand it can sell those Yen at a pre defined dollar rate.

Example

Macomb sold/exported products worth 10000 Japanese Yen. This is expected to be received after 1 month. To hedge Macomb also buys a Yen Put option to sell Yen at rate of $1 = 100 Japanese Yen.

Suppose, currently

$1 = 100 Japanese Yen.

So, Macomb is expected to get: 10000 / 100 Japanese Yen = $ 100

Now after 1 month exchange rate changes and Yen weaken

$1 = 110 Japanese Yen.

So, Macomb is expected to get: 10000 / 110 Japanese Yen = $ 90.909. But now Macomb has option.

It will use that option to get $1 = 100 Japanese Yen..

Hence, after 1 month also,

Macomb is expected to get: 10000 / 100 Japanese Yen = $ 100